Circular No. 9074

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 S1 (GIBBS) A. R. Gibbs reports his discovery of a comet on unfiltered CCD images

taken with the Catalina 0.68-m Schmidt telescope (discovery observation tabulated below); four co-added 30-s exposures show an 8′′ × 10′′ compact coma elongated in the direction of the faint 20′′ tail (p.a. 270^{o}), while four co-added 80-s frames from Sept. 20.43–20.46 UT show the narrow tail to be 45′′ long. After posting on the ‘NEOCP’ webpage, other CCD astrometrists have commented on the object’s cometary appearance. E. Guido, Castellammare di Stabia, Italy, writes that fourteen co-added unfiltered 120-s exposures taken by G. Sostero and himself remotely using a 0.25-m reflector near Mayhill, NM, on Sept. 21.4 show a condensed coma nearly 8′′ in diameter and a broad tail about 20′′ long in p.a. 260^{o}. H. Sato (Ota-ku, Tokyo, Japan, also remotely using a 0.25-m reflector near Mayhill on Sept. 21.4) reports a coma diameter of 12′′, slightly elongated in p.a. 320^{o}. R. Miles (2.0-m “Faulkes Telescope North” at Haleakala, Sept. 21.6) measures a 5′′ tail in p.a. 277^{o}. L. Buzzi (Varese, Italy, 0.60-m reflector, Sept. 22.1) notes that stacked images totalling 20 min exposure time show a strong central condensation and a faint tail ∼ 20′′ long in p.a. ∼ 270^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Sept.20.38191 3^{h}59^{m}54.^{s}60 −10^{o}17′18′′.3 18.6 Gibbs

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2009-S70.

T = 2009 July 24.990 TT  ω = 221.^{o}712
e = 0.35021  Ω = 155.805 }2000.0
q = 2.36593 AU  i = 25.378

a = 3.64108 AU n^{o} = 0.141860 P = 6.95 years

2007 TY_{430}

K. S. Noll, Space Telescope Science Institute (STSI); W. M. Grundy, Lowell Observatory; S. D. Benecchi, STSI; and H. A. Levison, Southwest Research Institute, report observations of the known transneptunian minorplanet binary 2007 TY_{430} (cf. MPEC 2008-M38; IAUC 8962) starting at 2008 Sept. 20.5253 UT with the Wide Field Planetary Camera 2 on the Hubble Space Telescope. At the time of these observations, the binary was separated by 0′′.248 ± 0′′.001 in α and 0′′.399 ± 0′′.001 in δ, with a difference between the two components of 0.22 ± 0.10 magnitude.

2009 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green